Friday, July 17, 2015

P390 continued DOS sysgen, Pertec disk drive restoration continues


I found a manual from a newer version of DOS (VSE) online which mentioned the error message 0I68A that I am receiving when my new supervisor fails to IPL. Based on that manual, the error code is telling me that I over-allocated the real storage of the Hercules simulated processor, putting too much of it into the real area thus not enough is there for the supervisor to complete IPL.

I had a bit of time at lunch and reassembled the supervisor, then booted it up to see how I did. I had a successful IPL of my system, without any of the customizing ADD and DEL commands previously required - simple IPL commands of SET and DPD to start up a DOS system.

Next up, I will play with the standard assignments, labels and procedures to get this where I want, generate my own Power/VS spooling system, then ensure it all comes up properly and works as intended. I started writing the JCL, macros and utility commands I needed tonight, to be as prepared as I could for when I 'powered on' the mainframe to test. That is the approach used with real systems of that era.


I had to replace some foam that sealed up the area where the heads enter the disk cartridge, since the original foam had become crumbling dust, exactly the opposite of what you want anywhere near the ultra thin spacing of the heads over a spinning disk. I had some suitable supply on hand and installed it today.

Since the timer board is not doing its job, I needed to have a schematic but none is available. Fortunately, it is a low density PCB with big visible traces that are easily followed. I whipped up a circuit diagram in the remainder of my lunch hour and in the late afternoon. Once done, I shared it back for those who might also need this documentation.

I started tracing a few voltages to figure out what was happening with the timer board function, as well as the solenoid lock and ready status. I see that the solenoid driver is behaving as I would expect it to, if transistor Q42 were an open circuit. It is a two transistor cascade where the first (Q43) pulls a small current through the solenoid that biases the higher power transistor Q42 on which increases the current.

My symptom is that the solenoid barely moves or has to be pushed. Measuring the voltage drop and absolute levels  shows me that the solenoid is getting about 5V across it, but should be dropping 25. The circuit values would cause this if Q42 didn't conduct and only the first stage was acting on the solenoid.  I will pull it out of the circuit and check it out.

The timer board senses the voltage from the solenoid to determine that it is in safe mode and should start the timer for the 5.5 minute power-down. Since it doesn't see the condition, its behavior is understandable.

I ran out of time tonight but my next steps will be to step through parts of the circuit, validating the expected voltages or signal states. I might have several other parts that failed due to the power anomaly that released the magic white smoke.


A number of people with 026 and 029 keypunches were looking for a source for the v-belt that is used inside the mechanism. I used one of the NOS belts from CHM to measure and study the part, then located some alternatives that should be close enough. Two are on the way to me for testing.

No comments:

Post a Comment